OpenStudy (anonymous):

Find the derivative of y=3tan^2 (x/3)

OpenStudy (solomonzelman):

This is what I found on yahoo. y = tan^2(x^3) y'(incomplete) = 2 * tan(x^3) Take care of trig function next. y'(incomplete) = 2 * tan(x^3) * sec^2(x^3) Take care of power function inside the trig function y' = 2 * tan(x^3) * sec^2(x^3) * 3x^2 = 6x^2 * tan(x^3) * sec^2(x^3) More elegant way? f(x) = x^2, g(x) = tan(x), h(x) = x^3 y = f(g(h(x))) = tan^2(x^3) Using chain rule y' = f'(g(h(x)) * g'(h(x)) * h'(x) f(x) = x^2, g(x) = tan(x), h(x) = x^3 f'(x) = 2x, g'(x) = sec^2(x), h'(x) = 3x^2 Substituting, y' = f'(g(h(x)) * g'(h(x)) * h'(x) = 2(tan(x^3)) * sec^2(x^3) * 3x^2 = 6x^2 * tan(x^3) * sec^2(x^3)

OpenStudy (anonymous):

yea but my equation is different

OpenStudy (solomonzelman):

can you derive each side of the problem?

OpenStudy (anonymous):

i know you bring in the 2 to the font so it will be y=6tan(x/3) but idk what to do next

OpenStudy (phi):

if the problem was y= 3x^2 you would do dy = 6x dx in this case your "x" is tan(x^3) so you would have dy = 6tan(x^3) d tan(x^3) in other words you know have to take the derivative of tan(x^3)

OpenStudy (anonymous):

how do i do that ?

OpenStudy (anonymous):

sec^2x?

OpenStudy (phi):

sec^2(stuff inside) you started with tan(x^3) so you would get sec^2(x^3) d x^3 you now take the derivative of x^3

OpenStudy (anonymous):

what happen to the 6?

OpenStudy (phi):

all the other stuff is still there. I was just concentrating on the derivative part. btw, is this the problem y=3tan^2 (x/3) ? (somehow x/3 turned into x^3) to summarize, d (y=3tan^2 (x/3) ) dy = 3 d tan^2 (x/3) dy = 6 tan(x/3) d tan(x/3) dy = 6 tan(x/3) sec^2(x/3) d x/3 dy = 6 tan(x/3) sec^2(x/3) ⅓ dx dy = 3 tan(x/3) sec^2(x/3) dx dy/dx = 2 tan(x/3) sec^2(x/3)

OpenStudy (phi):

to summarize, d (y=3tan^2 (x/3) ) dy = 3 d tan^2 (x/3) dy = 6 tan(x/3) d tan(x/3) dy = 6 tan(x/3) sec^2(x/3) d x/3 dy = 6 tan(x/3) sec^2(x/3) ⅓ dx dy = 2 tan(x/3) sec^2(x/3) dx dy/dx = 2 tan(x/3) sec^2(x/3)

OpenStudy (anonymous):

is that it?

OpenStudy (phi):

can you following the steps?

OpenStudy (anonymous):

yes kinda

OpenStudy (anonymous):

thank you

OpenStudy (phi):

does the answer match any of your choices?

OpenStudy (anonymous):

yes!:) thnxs